HUBBLE SPACE TELESCOPE DAILY REPORT #2163 PERIOD COVERED: 0000Z (UTC) 07/10/98 - 0000Z (UTC) 07/13/98 Daily Status Report as of 194/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Seven Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in 2.1 and in HSTAR 6580, during the fourth iteration of this proposal, the guide star acquisition failed to gyro control at 192/015955Z, and the take data flag went down for three observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.3 Completed Two Sets of NIC/1/2 7852 (Weighing the Stellar Halo of NGC 5907 with NICMOS) The Near Infrared Camera (Cameras One and Two) was used to make observations of the staler halo around NGC-5907 in order to determine its composition. This stellar halo has a radial profile that roughly traces the isothermal profile required to reproduce the rotation curve. As described in 2.1 and in HSTAR 6580, during the second iteration of this proposal, the guide acquisition failed to gyro control at 192/015955Z, and the take data flag went down for two observations. Subsequent reacquisitions during this proposal were successful. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.4 Completed Eleven Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.5 Completed Two Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. As described in 2.1 and in HSTAR 6580, during the second iteration of this proposal, the guide acquisition failed to gyro control at 192/015955Z, and the take data flag went down for two observations. Subsequent reacquisitions during this proposal were successful. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.6 Completed Two Sets of NIC/1/2/3 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the ultraluminous infrared galaxies QDOTULIR18 and QDOTULIR13. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. The observations were completed as planned, and no anomalies were noted. 1.7 Completed Three Sets of STIS/CCD/MA1 7288 (High Resolution Ultraviolet Spectroscopy of Hercules X-1/HZ Herculis) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make ultraviolet spectrographic observations of the X-ray binary pulsar 4U1656+35 (a.k.a., Hercules X-1/HZ Herculis). A wave length calibration observations was also taken. At the beginning of the second iteration of this proposal, the MCE-1 reset described in HSTAR 6582 and in 2.3 occurred, and ten STIS images were lost. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.8 Completed NIC/1/2 7821 (SINS: The Supernova Intensive Study - Cycle 7+: Scarlet SINS) The Near Infrared Camera (Camera One and Two) was used to make observations of the super nova LMC-SN1987 which is located in the Large Magellanic Cloud. Supernovae create the chemical history of the Universe, energize the interstellar gas, and stand at the center of efforts to understand the extragalactic distance scale. The observations were completed as scheduled, and no problems were noted. 1.9 Completed STIS/CCD 7476 (Sakurai's Nova-Like Object: Real-Time Monitoring of a Stellar Thermal Pulse) The Space Telescope Imaging Spectragraph (CCD) was used to make images of Sakurai's nova-like object in order to do real-time monitoring of a stellar thermal pulse. The eruption of Sakurai's nova-like variable in Sagittarius presents a "once-in-a-lifetime" opportunity for real-time observations of a star undergoing a final helium thermal pulse. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Two Sets of NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Nine Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.12 Completed NIC/2 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera Two) was used to make narrow band observations of the BL-Lac (BAL) QSO Q1439+0047. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 1.13 Completed Two Sets of WF/PC-2 6776 (A Comparative HST Imaging Study of Radio Galaxies and Hosts) A clear understanding of the nature of the host galaxies of radio-quiet quasars, radio-loud quasars and radio galaxies is of central importance in any attempt to unify and relate the three main classes of powerful active galaxies. Existing HST studies of quasar hosts, whilst producing many interesting new results, have yielded surprisingly little hard information on the properties of the hosts themselves. This imaging study of galaxy 3C79 using the WF/PC-2 will remedy this situation by undertaking redder, line-free observations of our three matched AGN samples. The proposal completed with no reported problems. 1.14 Completed NIC/2/3 7453 (The SBF Hubble Diagram) The Near Infrared Camera (Camera Two and Three) was used to make deep NICMOS observations of galactic cluster Abell-3565 to determine its distance using the surface brightness fluctuation {SBF} method. HST is essential for its high spatial resolution which makes it possible to detect and remove globular clusters and background galaxies that contaminate the SBF measurement. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 6837 (The Ultraviolet Properties of "Dwarf" Seyfert 1 Nuclei) The WF/PC-2 was used to make ultraviolet observations of the "dwarf", type one, Seyfert galaxy NGC-4501. This class of "dwarf" Seyfert populate the faint end of the local AGN luminosity function and hold important clues to the evolution of the entire AGN population. The observations were completed as scheduled, and no problems were reported. 1.16 Completed WF/PC-2 7385 (Low-Mass Halo Binaries) The WF/PC-2 was used to make observations of the low mass, halo, binary, {Population II}, M-type, sub-dwarf, star LHS364. This star is too faint to be resolved through ground-based speckle interferometry, thus HST represents the only means of probing separations. The observations were completed as planned, and no anomalies were reported. 1.17 Completed WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the rich star cluster NGC-1818 which is located within the Large Magellanic Cloud in order to study its formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.18 Completed NIC/1/2/3 7849 (Metallicity and the Cepheid Distance Scale) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of Cepheid variable stars in the NGC-4536 galaxy. The local distance scale to galaxies is largely based on the Cepheid period-luminosity relation. Motivated by the goal of measuring an accurate value of the Hubble constant, HST has been used successfully to discover scores of Cepheid variables in each of about 20 galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.19 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local Group galaxies offer the opportunity for a closer look at young stellar populations. The observations were completed as planned, and no anomalies were reported. 1.20 Completed NIC/1/2 7861 (A Search for Extrasolar Giant Planets in the Nearby TW Hya Association) The Near Infrared Camera (Cameras One and Two) was used to search for extrasolar giant planets on the nearby star CD-337795 (a.k.a., TW Hya). One of the primary goals of NASA's mission is the identification of formed and nascent planetary systems. The observations were executed as scheduled, and no anomalies were noted. 1.21 Completed FGS 7679 (Intermediate Plate Scale Calibration (FGS-3)) The Fine Guidance Sensors (FGS-3)was used to make observations of star HD65949 and several other reference stars in order to completed an intermediate plate scale calibration. The essential point of this calibration is to refine our knowledge of the plate scale in the central region of the astrometer FGS for this is where all annual parallax work will be executed by the GOs and GTOs. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 26 Successful : 25 HSTAR 6580 documents a guide star acquisition failure to gyro control at 192/015955Z when the coarse angle check failed. The attitude was good. Subsequent re-acquisitions were completely successful. The proposals described in 1.2, 1.3, and 1.5 were affected. Scheduled Reacquisitions : 23 Successful : 23 2.2 FHST Updates: Scheduled : 46 Successful : 46 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset ten times. Per ROP SR-3, the SSR SCI pointer was reset to recover science data for the DCF at 191/1751Z. ESTR-2 reconditioning was performed starting at 191/1943 per operations request and ROP RD-7. A TTR was written for a 15-minute loss of forward service at 192/0042Z due to equipment problems at STGT. Science operations were not affected. Another TTR was written to document a 35-minute period of unrecoverable realtime engineering data and a 4-minute, 29-second period of recoverable realtime data. This event started at 192/0442Z. The STIS MCE-1 reset at 192/133941Z while the MAMA-1 high voltage was on and while outside an SAA interval. The proposal described in 1.7 was affected. The STIS error counter was reset at 192/1437Z via ROP NS-12. Recovery occurred via the SMS at 193/1201Z, the next scheduled time for the MAMA-1 high voltage turn on. For all three days of the weekend, the DF-224 SPC memory were dumped to CCS strings A and B. Both were successful the first day; for the second day B string was down; and for the third day there were problems on both strings. Later, this third day's dump was successful to A string only. At 193/0716Z, a full DF-224 memory dump was successfully performed, sending data to the operational string and to CCS string A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH